OpenStudy (debpriya):

if the product of all the factors of a number N is equal to the square of the number and the sum of all the factors of the number other than the number itself is 21, then find the number of possible values for the number? a. 0 b.1 c.2 d. more than 2

ganeshie8 (ganeshie8):

\(n^{\tau(n)/2} = \prod\limits_{d\mid n}d = n^2\) That means \(\tau(n) = 4\)

OpenStudy (debpriya):

yes yes !

ganeshie8 (ganeshie8):

I'm using n = N

ganeshie8 (ganeshie8):

You familiar with the number theoretic functions \(\tau\) and \(\sigma\) ?

ganeshie8 (ganeshie8):

\(\tau(n)\) is number of positive factors of \(n\) \(\sigma(n)\) is sum of positive factors of \(n\)

OpenStudy (debpriya):

yep so the way we can write the number will be p^1*^1 and p^3 right ? based on the previous question ?

OpenStudy (debpriya):

i meant p^1 * q^1 where p and q are prime

ganeshie8 (ganeshie8):

Excellent! only below forms are possible for \(n\) to have exactly 4 positive factors : \(p^3\) \(p^1q^1\)

ganeshie8 (ganeshie8):

What are the factors when of \(n\) when \(n=p^3\) ?

OpenStudy (debpriya):

the factors will be only p right ?

ganeshie8 (ganeshie8):

factors of \(n\) are the integers that divide \(n\)

OpenStudy (debpriya):

oh and p square ?

ganeshie8 (ganeshie8):

think a bit, there should be exactly \(4\) positive factors of \(p^3\) right ?

OpenStudy (debpriya):

so the factors are 1,p, p square and p cube ?

ganeshie8 (ganeshie8):

Yes!

ganeshie8 (ganeshie8):

what are the factors of \(p^1q^1\) ?

OpenStudy (debpriya):

1,p,q and pq

ganeshie8 (ganeshie8):

so we need to solve 1 + p + p^2 + p^3 = 21

ganeshie8 (ganeshie8):

and independently 1 + p + q + pq = 21

ganeshie8 (ganeshie8):

Key things to keep in mind to arrive at the answer quick : 1) 21 is an odd integer 2) sum of four odd integers is always even 3) 2 is the only even prime number

OpenStudy (yttrium):

What subject is this?

OpenStudy (debpriya):

there fore p+q+pq= even , thus p ,q and pq should also be even. am i right?

ganeshie8 (ganeshie8):

Also keep in mind p and q are two "different" primes

ganeshie8 (ganeshie8):

number theory @Yttrium

ganeshie8 (ganeshie8):

Do we have two different even primes ?

OpenStudy (debpriya):

nope

OpenStudy (debpriya):

so this wont work :(

ganeshie8 (ganeshie8):

so 1+p+q+pq = 21 has no solutions

ganeshie8 (ganeshie8):

How about the other equation 1 + p + p^2 + p^3 = 21 ?

OpenStudy (debpriya):

here p+p^2+p^3 should be even ....so in this case p=2

ganeshie8 (ganeshie8):

Yes, plugin p=2 and see if it satisfies the above equation

OpenStudy (debpriya):

it doesn't satisfy the equation

ganeshie8 (ganeshie8):

can we conclude there exists no integer \(n\) that satisfies the given conditions ?

OpenStudy (debpriya):

so answer is 0?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (debpriya):

i don't know why but in the solution set they have marked it as 2. But this method did make sense

ganeshie8 (ganeshie8):

Oh, we made a mistake

ganeshie8 (ganeshie8):

the sum of all the factors of the number `other than the number itself` is 21

ganeshie8 (ganeshie8):

we should exclude the number itself from the sum, but we have added all the factors

ganeshie8 (ganeshie8):

we should be solving the equations : 1 + p + p^2 = 21 and 1 + p + q = 21

ganeshie8 (ganeshie8):

try again

OpenStudy (debpriya):

in the first equation p has to be odd..coz 2 doesn't satisfy...

OpenStudy (debpriya):

but the first equation has no such odd number only the second equation can we satisfied by taking two odd prime number below 20 which are 3,17 and 7 ,13 therefore the answer is 2..am i right?

ganeshie8 (ganeshie8):

Sounds good to me!

OpenStudy (debpriya):

thank you once again ! You have been a great help :)

ganeshie8 (ganeshie8):

Np:)